the measure of an angle is 12 degrees less than twice the measure of its suppliment. what is the measure of this angle?
Answer by stanbon(75887)

the measure of an angle is 12 degrees less than twice the measure of its supplement. what is the measure of this angle?
Let the supplement be "x":
Then angle = 2x-12
---
Equation:
x + 2x-12 = 180
3x = 192
x = 64 degrees
The angle = 2*64-12 = 116 degrees
====================
Cheers,
Stan H.